OpenStudy (anonymous):

Use the Intermediate Value Theorem to prove that the function f (x) = x^2-4x+3 intersects the x-axis on the interval [0,2]. Can we use the same theorem to say the same about the function g(x) =x^2-2x+1/x-1? ANSWER EITHER YES OR NO ? Not sure how to go about doing this ? do you fill in 0,1... and so on for f(x) until you get a positive and Negative ?

OpenStudy (anonymous):

 if the graph passes x-axis it will have a zero in the interval and on both sides of zero the sign of function will change check f(0) and f(2) if the sign changed means there is a zero in there

OpenStudy (anonymous):

yes i did that previous and the signs changed which means a negative by a positive is a negative therefore its continuous . but came we same the same for the second part or how do you go about solving it ??? answer only allows me to say yes or no ??

OpenStudy (anonymous):

you can not use the intermediate value therom on the second part becuase it is not continuous on the interval from [0.2]

OpenStudy (anonymous):

ok cheers

OpenStudy (anonymous):

f(x) is continuous for all x, hence it is continuous on interval [0,2]. So we can apply the IVT -1 = f(2) < 0 < f(0) = 3 therefore there must be a c in [0,2] such that f(c)=0 g(x) =(x^2-2x+1)/(x-1) is not continuous when x=1 so the IVT doesn't apply over any interval containing x=1